OpenStudy (anonymous):

The sum of a number and 6 times its reciprocal is 7. Find the number. If the problem has two solutions, write your answers in the box below.

OpenStudy (anonymous):

x + 6/x = 7 x + 6 = 7x x - 7x = -6 -6x = -6 x = 1

OpenStudy (anonymous):

So it's -6,1 right? :P

OpenStudy (anonymous):

the answer is 1.

OpenStudy (anonymous):

Oh okay thanks